# Otter-Knowledge: benchmarks of multimodal knowledge graph representation learning from different sources for drug discovery

### v1

@@ A.3 Datasets release
We release the 4 KGs used to create the 12 pretrained models shown in (Table 7). The KGs were created from the original dataset sources using a JSON schema that specifies the type of entities and their namespaces; we use namespaces to assign a unique identifier to each entity based on its source, so that we can connect them within and across dataset sources (an exception is PrimeKG, which requires custom processing of nodes because of the format of the original CSV files - we use however the same relations as defined by PrimeKG). For each entity type, the schema defines the relevant object and data properties and the modalities for both entities and attributes.
@@
Figure 4 shows an example schema for the UBC KG, which comprises of 9,386,418 RDF triples obtained by combining factual information from Uniprot, BindingDB and a subset of ChemBL Drugs molecules obtained from the Open Targets platform [19]; however for training the GNN we omitted triples with the rdf:type predicate, as this information is implicit in the modality and it would result in a very densely connected graph (e.g, when considering all nodes of type Protein), resulting in 6,207,654 triples. Figure 5 shows a portion of the resulting KG where an entity of type Protein defined with its UniProt ID and namedspace (https://www.uniprot.org/uniprot/P06820) has data attributes such as sequence, label (Neuraminidase), family (glycosyl hydrolase 34 family), as well as object properties from and to other entities (e.g, a target_of relation to drug https://go.drugbank.com/drugs/DB02268, which is also the object of a owl:sameAs relation from https://www.ebi.ac.uk/chembl/target_report_card/CHEMBL1091644 ). We note that for STITCH the raw data was additionally processed to filter out interactions with the highest likelihood (> 0.9) and to remove duplicated interactions based on chemical SMILES and protein sequence. Also, files containing information about interactions, SMILES and protein sequences were joined into a single CSV file for which schema was designed for KG creation.
Figure 4: Schema for UBC Knowledge Graph.
@@
The datasets are serialised in N-Triples  [^10] https://www.w3.org/TR/n-triples/, a line-based, plain text serialisation format for RDF triples and the pretrained models are accessible at  [^11] https://github.com/IBM/otter-knowledge. The original tabular datasets were either downloaded from their corresponding websites or accessed using IBM Deep Search  [^12] https://ds4sd.github.io, such as in the case of Uniprot SwissProt. Leveraging state-of-the-art AI methods, Deep Search collects, converts, enriches, and links large document collections [32] and it comes pre-loaded with millions of standard technical documents from many public data sources.
@@
The same way as the models, the datasets can be found in the Huggingface hub. The names of the datasets in the hub, under the IBM organization space, are the next ones:
  • otter_uniprot_bindingdb: (UB) MKG comprising entities and relations from Uniprot and BindingDB.
@@
  • otter_uniprot_bindingdb_chembl: (UBC) MKG comprising entities and relations from Uniprot, BindingDB and ChemBL.
  • otter_dude: (DUDe) MKG with data from the preprocessed version of DUDe
  • otter_primekg: (PrimeKG) MKG with data from the PrimeKG
  • otter_stitch: (STITCH) MKG with data from the preprocessed version of STITCH
@@
Figure 5: Subset of a KG in UBC. Attribute nodes are coloured in green, type classes in blue, uniprot, chembl, drugbank and bindingDB entities are coloured in orange, salmon, blue and purple respectively
A.4 Results with UDB and UB
The outcomes of the model that underwent pretraining on Uniprot and BindingDB (UB) compared to the model that underwent pretraining on Uniprot, Drugbank, and BindingDB (UDB) are presented in Table 8. Although we are unable to share the processed Drugbank datasets because of licensing limitations, it is evident that incorporating Drugbank does not lead to significantly different results compared to using UB alone.
@@ A.4 Results with UDB and UB
Table 8: Results of knowledge enhanced representation on three standard drug-target binding affinity prediction benchmarks datasets with different splits of the models pretrained with UDB and UB respectively.
Datasets (UDB) DTI DG DAVIS KIBA
@@
Table 9: Universality results for UDB, where the scoring function considers all the knowledge graph links during the training. The table results should be compared with the results in Table 3 (UDB). The evaluation metrics is Pearson correlation (higher is better).
A.5 Computing resource and training time
@@
For managing the large graphs like the UBC, we utilized 200GB of CPU memory. The storage of the initial embeddings for the modality required the highest amount of memory compared to other components. With the implementation of the GAS methods explained in Section 2.3, the GNN training process was able to handle the large graphs effectively, even when constrained by GPU memory limitations.
A.6 Downstream task open source
As we have discussed in Subsection 4.1, the neural architectures in the TDC framework are not optimal for the TDC DG benchmarks because they are designed to learn proteins/drugs representation from scratch. For pretrained proteins/drugs representation evaluation we created a new architecture that provide better results for both the baseline and the GNN representation. We made these architectures available for reproducible results. Additionally, we have open-sourced the source code of the ensemble methods used in our work.

### v3

@@ A.3 Datasets release
Figure 4: Schema for Otter-UBC Knowledge Graph.
Figure 5: Subset of a KG in Otter-UBC. Attribute nodes are coloured in green, type classes in blue, uniprot, chembl, drugbank and bindingDB entities are coloured in orange, salmon, blue and purple respectively
@@
We release the 4 KGs used to create the 12 pretrained models released. These KGs were constructed from the original dataset sources using a JSON schema that specifies the type of entities and their respective namespaces. We use namespaces to assign a unique identifier to each entity based on its source, so that in the KGs we can connect and cross reference entities within and across dataset sources (an exception is PrimeKG, which requires custom processing of nodes because of the format of the original CSV files - we use however the same relations as defined by PrimeKG). For each entity type, the schema defines the relevant object and data properties and the modalities for both entities and attributes.
Duplicates are therefore handled effectively through the user provided schema by specifying the namespace of the original source and the unique entity id (e.g, if an entity in BindingDB references an entity with a Uniprot ID, the schema needs to specify that the target entity belongs to the Uniprot namespace). Additionally, an owl:sameAs relation can be specified in the schema to leverage the links explicitly established across entities with different IDs across data sources. Nonetheless, coping with the intricacies of automatic alignment for entities not explicitly link in the original sources, along with relations that may be semantically similar yet have different names across datasets, constitutes a research challenge. The down effect of this is that the model may face difficulties to learn to transfer the commonalities between two (duplicated) entities or relations, if there isn’t an explicit mechanism to force the model to do so. Designing a method that enables the model to learn to transfer across sources with varying schemas remains an open research question.
@@
Figure 4 shows an example schema for the Otter-UBC KG, which comprises 9,386,418 RDF triples obtained by combining factual information from Uniprot, BindingDB and a subset of ChemBL Drugs molecules obtained from the Open Targets platform [19]; however for training the GNN we omitted triples with the rdf:type predicate, as this information is implicit in the modality and it would result in a very densely connected graph (e.g, when considering all nodes of type Protein), resulting in 6,207,654 triples.
Figure 5 shows a portion of the resulting KG where an entity of type Protein defined with its UniProt ID and namedspace (P06820) has data attributes such as sequence, label (Neuraminidase), family (glycosyl hydrolase 34 family), as well as object properties from and to other entities (e.g, a target_of relation to drug DB02268, which is also the object of a owl:sameAs relation from CHEMBL1091644 ).
We note that for STITCH the raw data was additionally processed to filter out interactions with the highest likelihood (greater than 0.9) and to remove duplicated interactions based on chemical SMILES and protein sequence. Also, files containing information about interactions, SMILES and protein sequences were joined into a single CSV file for which schema was designed for KG creation.
@@
The datasets are serialised in N-Triples [^9] https://www.w3.org/TR/n-triples/, a line-based, plain text serialisation format for RDF triples and the pretrained models are accessible at the link  [^10] https://github.com/IBM/otter-knowledge. The original tabular datasets were either downloaded from their corresponding websites or accessed using Deep Search [^11] https://ds4sd.github.io, such as in the case of Uniprot SwissProt. Leveraging state-of-the-art AI methods, Deep Search collects, converts, enriches, and links large document collections [32] and it comes pre-loaded with millions of standard technical documents from many public data sources.
@@
The same way as the models, the datasets have been released and can be found in the Huggingface hub. The names of the datasets in the hub, under the IBM organization space, are the next ones:
  • otter_uniprot_bindingdb: MKG comprising entities and relations from Uniprot and BindingDB.
@@
  • otter_uniprot_bindingdb_chembl: (Otter-UBC) MKG comprising entities and relations from Uniprot, BindingDB and ChemBL.
  • otter_dude: (Otter-DUDe) MKG with data from the preprocessed version of DUDe
  • otter_primekg: (Otter-PrimeKG) MKG with data from the PrimeKG
  • otter_stitch: (Otter-STITCH) MKG with data from the preprocessed version of STITCH
@@
A.4 Information control flow and Restricted (not noisy) link prediction
The graph might be large, with many links that may not be relevant for the tasks ahead. In the process of GNN pretraining , substantial amounts of known attributes are available for both proteins and molecules in the KGs. However, when it comes to fine-tuning for specific tasks like predicting drug-target binding affinity, this information is severely limited for novel proteins and drugs. In such cases, only basic details such as protein amino acid sequences and molecule SMILES are accessible. We aim to study if this discrepancy in data between pretraining and fine-tuning phases could lead to a decrease in the accuracy of binding affinity prediction for newly introduced entities.
To assess the impact of i) regulating the information propagation / flow and ii) restricted link predictions, we have a list of permissible specific links in each KG that matter for predicting drug-target binding in downstream tasks. Our tests showed results for both cases: when we control the information flow and predict links only from our picks, and another where we considered all links regardless of the task.
In the information control flow process within each iteration of GNN training, exclusively the messages originating from approved links are employed to update node embedding. As illustration, consider a protein node; the accepted link types to control the information flow include sequence attributes. Similarly, for a molecule node, permissible link types encompass the SMILES attributes. This selection stems from the fact that, during the testing phase, newly introduced Proteins and Molecules lack attributes beyond sequence and SMILES.
For the experiments with restricted links prediction, these are the ones we chose for each knowledge graph:
  • Otter-UBC: target of, binding to
  • Otter-DUDe: binding to
  • Otter-PrimeKG: protein protein interaction, drug drug synergistic-interaction, drug protein transporter, drug protein target, drug protein enzyme
  • Otter-STITCH: binding to
The experimental results without control flow and link predictions with all possible links, including potentially noisy links to the downstream tasks, are demonstrated in Table 4. The results prove our method works well even with noisy links, as the outcomes are comparable in both cases – whether we limit the links or not. Additionally, using extra modalities in pre-training do not impair generalisation to fine-tuning.
A.5 Results with Otter-UDB and Otter-UB
The outcomes of the model that underwent pretraining on Uniprot and BindingDB (Otter-UB) compared to the model that underwent pretraining on Uniprot, Drugbank, and BindingDB (Otter-UDB) are presented in Table 8. Although we are unable to share the processed Drugbank datasets because of licensing limitations, it is evident that incorporating Drugbank does not lead to significantly different results compared to using UB alone.
@@ A.5 Results with Otter-UDB and Otter-UB
Table 8: Results of knowledge enhanced representation on three standard drug-target binding affinity prediction benchmarks datasets with different splits of the models pretrained with Otter-UDB and Otter-UB respectively.
Datasets (Otter-UDB) DTI DG DAVIS KIBA
@@
Table 9: Universality results for Otter-UDB, where the scoring function considers all the knowledge graph links during the training. The table results should be compared with the results in Table 3 (Otter-UDB). The evaluation metrics is Pearson correlation (higher is better).
A.6 Computing resource and training time
@@
For managing the large graphs like the Otter-UBC, we utilized 200GB of CPU memory. The storage of the initial embeddings for the modality required the highest amount of memory compared to other components. With the implementation of the GAS methods explained in Section 2.3, the GNN training process was able to handle the large graphs effectively, even when constrained by GPU memory limitations.
A.7 Downstream task open source
As we have discussed above, the neural architectures in the TDC framework are not optimal for the TDC DG benchmarks because they are designed to learn proteins/drugs representation from scratch. For pretrained proteins/drugs representation evaluation we created a new architecture that provide better results for both the baseline and the GNN representation. We made these architectures available for reproducible results. Additionally, we have open-sourced the source code of the ensemble methods used in our work.
A.8 Results with Spearman rank correlation
In this section, we present supplementary experiments involving the Spearman rank correlation, an alternative metric to Pearson’s correlation.
The outcomes presented in Table 10 were collected from the model pretraining on UBC knowledge graphs. Evidently, the augmented representation utilizing data from UBC knowledge graphs outperforms the outcomes obtained from utilizing basic pretrained embeddings from ESM models for proteins and the Morgan Fingerprint for drugs. These findings underscore the robustness of the enhanced representation across diverse evaluation metrics.
Datasets DTI DG DAVIS KIBA
Splits Temporal Random Target Drug Random Target Drug
Otter Distmult 0.590 0.635 0.510 0.146 0.837 0.570 0.549
Otter TransE 0.592 0.638 0.506 0.128 0.833 0.599 0.536
Otter Classifier 0.593 0.637 0.512 0.102 0.840 0.580 0.579
ESM+Morgan Fingerprint 0.591 0.634 0.504 0.268 0.826 0.579 0.536
Table 10: The Spearman’s rank correlation of the models trained on Uniprot, BindingDB and ChemBL (UBC) and the baseline models using ESM and Morgan Fingerprint embeddings for proteins and drugs respectively.